OpenStudy (anonymous):

In the binary system, the number 145 is represented as a. 01010101 b. 01101100 c. 10010001 d. 11001001

OpenStudy (anonymous):

is it B?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

and its not a soo what is it?

myininaya (myininaya):

\[\lfloor{\frac{\log(145)}{\log(2)}} \rfloor=7\] \[1 \cdot 2^7+0 \cdot 2^6+ 0 \cdot 2^5+ 1 \cdot 2^4+0 \cdot 2^3+ 0 \cdot 2^2+ 0 \cdot 2^1 +1 \cdot 2^0\]

OpenStudy (zarkon):

I'm sure myininaya is going to tell you....there you go ;)

OpenStudy (anonymous):

so c or d?

myininaya (myininaya):

10010001

myininaya (myininaya):

the coefficients of 2^(whatevers) in the order i gave you

OpenStudy (anonymous):

ohhh okay thks

OpenStudy (zarkon):

I like converting it to hexadecimal...it's quicker... \[154_{10}=91_{16}\] then 9=1001 and 1=0001 that gives 10010001

myininaya (myininaya):

i think there is another way you can just do some divisions i believe